Ross Textiles wishes to measure its cost of common stock equity. The​ firm's stock is currently selling for ​$43.24. The firm just recently paid a dividend of

​$4.15. The firm has been increasing dividends regularly. Five years​ ago, the dividend was just $2.97. After under pricing and flotation​ costs, the firm expects to net

​$39.78 per share on a new issue.

a.  Determine average annual dividend growth rate over the past 5 years. Using that growth​ rate, what dividend would you expect the company to pay next​ year?

b.       Determine the net​ proceeds, Nn​, that the firm will actually receive.

c.  Using the​ constant-growth valuation​ model, determine the required return on the​ company's stock, r Subscript s, which should equal the cost of retained​ earnings,

r Subscript r.

d.  Using the​ constant-growth valuation​ model, determine the cost of new common​ stock, r Subscript n
Answer #1

Answer to Part a.

Current Dividend = $4.15
Dividend paid 5 years ago = $2.97
Time = 5 years

Future Value = Present Value * (1 +r)^ n
$4.15 = $2.97 * (1 + r)^ 5
(1 + r)^ 5 = 1.3973
1 + r = 1.3973 ^ (1/5)
1 + r = 1.06920
r = 0.06920
or r = 6.92%

Average Annual Dividend growth is 6.92%

Expected Dividend = Current Dividend * (1 + Growth Rate)
Expected Dividend = $4.15 * (1 + 0.0692)
Expected Dividend = $4.44

Expected dividend to pay next year is $4.44

Answer to Part b.

Net Proceeds = $39.78

Answer to Part c.

Required Return = Expected Dividend / Current Price + Growth Rate
Required Return = $4.44 / $43.24 + 0.0692
Required Return = 0.1027 + 0.0692
Required Return = 0.1719 or 17.19%

Answer to Part d.

Required Return = Expected Dividend / Current Price + Growth Rate
Required Return = $4.44 / $39.78 + 0.0692
Required Return = 0.1116 + 0.0692
Required Return = 0.1808 or 18.08%